# Balancing Confidentiality and Transparency for Blockchain-based Process-Aware Information Systems

### v2

@@ 4.2 Information viewpoint
Figure 3 depicts the configure functionality.
The process owner sends the process specification files (like the diagram in Fig. 1) to the Process Manager (1.0). The Process Manager, in turn, (1.1) processes the input data, storing the behavioral and business constraints it reports. Also, (1.2) it builds a policy template, associating roles of the process with the data they can access via encryption directives. Once created, the policy template is returned to the process owner.
In our running example, an extract of a policy template is:
{medicalPrescription:($PID and (PATIENT or RADIOLOGY)),
registration:
($PID and (RADIOLOGY or PATIENT)), report:($PID and (RADIOLOGY or WARD))}.
Here medicalPrescription, registration, and report represent messages, while PATIENT, RADIOLOGY and WARD are the roles of users who can access them. $PID is a placeholder for the process instance identifier (a piece of information that can be known only at runtime).
We give the process owner the option to revise the templates. In this way, adding custom roles that are not expected active parties in the process but need access to data. This is the case, for instance, of external auditors who need to certify data and its compliance with norms, laws and regulations. Table 1 shows an extract from the final access policies of our running example. In this case, each message is attached to a policy with an ID representing the process instance and the authorized role. Considering the medicalPrescription message, the policy specifies that to access it, a user has to be attested to have the role of a MINISTRY-INSPECTOR, or again to be a PATIENT or RADIOLOGY role within the running process instance.
@@
Once policies are confirmed, (1.3) they are sent to the Confidentiality Manager, which (1.4) stores them in the Distributed FS. Their permalink (1.5) is passed back to the Confidentiality Manager which saves it.
@@
Figure 4 depicts the instantiate functionality. Differently from the configure functionality, needed once per process, this one is designed for every new instantiation thereof. To initialize a new process instance (e.g., number 476948), participants send their data to the Process Manager (2.0) to authenticate themselves as participants in the process with their role (e.g., that user 0xc5…Fd43 operates as a WARD clerk in a new process instance PID476948). (2.1) The authentication is attested to by certifiers confirming the user’s self-declared roles.
@@
When all process participants are identified and associated with a role, (2.2) the Process Manager merges the received information with the process data acquired in step 1.1. Thereupon, (2.3) a selected participant kickstarts a new process instance through the Process Manager associated with the aforementioned information on the Process Contract, thereby initializing the instance’s public state. Notice that also this crucial step is notarized on chain (2.4). The selection of the participant, or participants, on behalf of whom the transaction is signed, transcends the scope of this paper and depends on the decision process and guidelines of the consortium.
@@
After initializing the process state, (2.5) the Process Manager triggers the Confidentiality Manager, which (2.6) stores the policy locator previously saved on the Confidentiality Contract. Finally, (2.7) the Confidentiality Manager notarizes the operation.
@@
The transact functionality reported in Fig. 5 represents the exchange of information during the execution of an activity in the process instance. This passage updates the public state update in the blockchain according to previous enforcement (e.g., activity to execute, expected participant, etc.) and manages accessibility in case of public and confidential cases.
The first case refers to public data and is displayed in Fig. 5(a).
@@
Initially, (3.0a) a user sends the execution data to the Process Manager in the context of an activity. For example, the WARD clerk identified by 0xc5…Fd43 replies to the
@@
“Check availability” activity request by the RADIOLOGY user within process instance 476948. To do so, the clerk sends a “Check availability” response message wherein the accepted variable is assigned with true since the appointment is confirmed alongside an available date. The boolean answer (accepted) is public data that is used later to make control-flow decisions and advance the state of the process, defining the next activity to perform. In our scenario, the exclusive choice gateway is based on the value of the accepted variable.
Then, (3.1a) the Process Manager invokes the Process Contract to request a state update, sending the execution data.
@@
In turn, (3.2a) the Process Contract enforces the business and control-flow logic through the run of its routine (based on previously agreed upon and immutable code, see step 1.1), which checks if the executed activity is enabled and if the message sender (in this case, user 0xc5…Fd43) is the expected one (a WARD), finally updating the state if the verification outcome is positive.
@@
Also, (3.3a) this step advances the state of the process, updating the expected next activity (here, “Confirm appointment”) and participant (RADIOLOGY). This piece of information is notarized into the Distributed Ledger.
@@
The confidential case is depicted in Fig. 5(b). The Process Manager, (3.0b) once it has received the data, (3.1b) forwards it to the Confidentiality Manager. The latter (3.2b) retrieves the policy from the Distributed FS, and (3.3b) uses it to encrypt the data via Attribute-Based Encryption (ABE). For example, considering Fig. 1 and Table 1, the medicalPrescription of the patient is encrypted with the following policy: {MINISTRY-INSPECTOR or (PID476948 and (PATIENT or RADIOLOGY))}.
@@
In particular, the Confidentiality Manager retrieves the specific authorization for the data to be stored and embeds it in the encryption mechanism, ensuring that only authorized participants can access it.
@@
Once encrypted, (3.4b) the Confidentiality Manager stores data in the Distributed FS, (3.5b) its locator in the Confidentiality Contract, and (3.6b) notarizes the operation in the Distributed Ledger.
@@
The execution of this routine by the Confidentiality Contract includes a call to the Process Contract (3.7b) to request a state update (enclosing the execution data). Following the example, this data is the activity being executed (“Check availability”), the user that performs it (e.g.,, PATIENT, along with their corresponding address 0x2e…6dd9), and the process instance (e.g.,476948).
Then, (3.8b) the Process Contract checks the current public state to update it. The operations involving the smart contracts (3.9b) are notarized in the Distributed Ledger.
The inspect functionality pertains to the access to message data. The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above). To read public data, a process participant or an auditor may send a request to the Process Manager, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester. This is the case, e.g., for the accepted field of the “Check availability” task. Reading confidential data requires more passages, illustrated in Fig. 6, and requires a preliminary step. All users should request at least once (and at every attribute update) a decryption key (Fig. 6(a)), to be used for every message they are going to read (Fig. 6(b)). For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Manager (4.0a). The latter (4.1a) forwards the request to the Confidentiality Contract, which is thus (4.2a) notarized on the Distributed Ledger. Then, (4.3a) the Confidentiality Manager generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via Attribute-Based Encryption (ABE), and (4.4a) sends it back. The decryption of confidential data is depicted in Fig. 6(b). Once the process participant has received the a-b key, they can use it to decrypt the confidential data. To do so, (4.0b) the user makes a request to the Confidentiality Manager. Then, (4.1b) the Confidentiality Manager retrieves the encrypted data from the Distributed FS and (4.2b) forwards it to the user. At this point, the latter (4.3b) uses the a-b key to decrypt the confidential data via ABE. Only if the user’s attributes satisfy the policy previously used to encrypt the data, can the user’s a-b key decrypt the confidential data (and the user inspect it in-clear).
@@
We remark that an a-b key stays with the user and lasts for as long as their attributes are not updated (in case, e.g., a user is involved in a new process instance, say 476949, a new attribute gets associated to them: $PID476949). The same key will be used to access all messages. For example, consider user 0xB0…1AA1 (who holds the attributes PID476948 and PATIENT).
As per the policies in Table 1, their a-b key can decrypt both the medical prescription and the registration alike, but not the report.
The a-b key of user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
@@
The key, in other words, is associated to the user and it is the sole instrument to access a document.
Therefore, the Confidentiality Manager does not need to generate a different decryption key for every document and give the same copy to multiple users, nor create multiple copies of the same document for different decryption keys (one per user), thus saving on unnecessary data replications and key distributions. Also, as the key stays with the users, and the documents are stored and notarized on tamper-proof external infrastructures (the Distributed FS and Blockchain Platform, respectively) the information remains available even should the CONFETTY system stop functioning. These aspects have an impact on security considerations, as we will discuss next.
@@ 5 Evaluation
To assess the feasibility of the CONFETTY architecture we implemented a supporting prototype and evaluated it. We first describe the implemented architecture analyzing its thread model, then we evaluate CONFETTY in operation.
@@ 5.1 Implementation
Figure 7 illustrates the technologies utilized in the implementation of CONFETTY. Our prototype is deployed on three main tiers. An application server hosts the Confidentiality Manager and the Process Manager, developed in Python 3.6.9 and Java SDK-13, respectively. Users communicate with it via TLS communication channels [37]. We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: archive.trufflesuite.com/ganache. Accessed: December 13, 2024. as the Distributed Tamper-proof File Storage. We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js.1 The source code of CONFETTY, alongside the experimental data we use, is publicly available at github.com/apwbs/CONFETTY.
5.2 Threat Model
In this section, we identify potential security threats that could compromise data confidentiality within our solution, and outline how the CONFETTY architecture addresses them. Our threat analysis is based on the STRIDE framework [35], a theoretical model that categorizes threats into six distinct groups: spoofing (i.e., impersonation of a legitimate entity), tampering (i.e., unauthorized modification of data to compromise its integrity), repudiation (i.e., the denial of having performed a particular action), information disclosure (i.e., the unauthorized exposure of sensitive data), denial of service (DoS, i.e., the disruption or degradation of system availability), and elevation of privileges (i.e., the unauthorized acquisition of higher-level access rights).
As said in Sect. 4.1, we assume users and components to be honest but curious and trusted, respectively. The possible threats thus reside in the communication between them, subject to potential attacks. Figure 8 displays a Data Flow Diagram (DFD), in which the orange dotted lines indicate the trust boundaries crossing data exchange points between different perimeters, represented by black arrows connecting different elements. The opening side of each boundary indicates the trusted component in the information exchange.
The CONFETTY Boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditor). In this case, a spoofing attack could result in a user impersonating another, but since we assume that each user is linked to the Blockchain Platform credentials, this attack is prevented. As for tampering, trust is based on the fairness of the user who sends the data. It cannot be the case for a user to repudiate an action because all of them are recorded on the blockchain with a transaction signed with the user’s private key. An attacker could sniff the data a user sends to CONFETTY. However, for those communications we resort to TLS channels, specifically designed to prevent information leakage [37]. DoS attacks are in principle feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform. Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1.0 in Sect. 4.2) which malicious nodes could exploit with multiple calls and bulky payloads. However, DoS attacks would lead to a disruption or degradation of the system availability, yet all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and from the Distributed Tamper-proof File Storage so the system can be reactivated or replicated without any data loss. Information disclosure is prevented by the fact that all sensitive information is encrypted via ABE. Elevating a user’s privileges is not possible since attributes and permissions are linked to a blockchain account.
The Blockchain Platform is the trusted component in the Blockchain Boundary (BB). In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the real one. This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their own private keys. Information disclosure can be achieved only on public data or resource locators of sensitive data. However, the latter is stored upon ABE-encryption, making its content unreadable by malicious users. Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols. As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack, and elevation of privileges, are counteracted by design, too. Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain. Since we resort to ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.

### v3

@@ 4.2 Information Viewpoint
The first functionality is configure. Here the process owner sends the process specification files (like the diagram in Fig. 1) to the Process Interface. The Process Interface, in turn, processes the input data, storing the behavioral and business constraints it reports. Also, it builds a set of parametric access policies, associating roles of the process with the data they can access via encryption directives (meeting Table 1). Once created, the policies are returned to the process owner. The one associated with the medical prescription (sent by the patient to the radiology clerk to take an appointment in Fig. 1), e.g., reads “$PID and (PATIENT or RADIOLOGY)”. Here, PATIENT and RADIOLOGY are the attributes representing the actors’ roles in the process. Not all patients and radiology clerks should access any medical prescription. Hence $PID, an attribute placeholder for the process instance identifier (a piece of information that can be known only at runtime), restricting access to the sole actors involved in that process instance. We give the process owner the option to revise the policies by adding custom roles that are not expected active parties in the process but need access to data (as per Table 1). This is the case, for instance, of external auditors who need to certify data and its compliance with norms, laws and regulations. Considering the medical prescription, the revised policy specifies that a user has to be attested to have the role of a MINISTRY-INSPECTOR, or again to be a PATIENT or RADIOLOGY within the running process instance: “MINISTRY-INSPECTOR or ($PID and (PATIENT or RADIOLOGY))”. Table 2 shows an extract from the final access policies of our running example. Once policies are confirmed, they are sent to the Confidentiality Interface, which stores them in the Distributed FS. Their permalink is passed back to the Confidentiality Interface to save it.
Figure 4 depicts the instantiate functionality.
@@
Differently from the configure, needed once for each process, this one is designed for every new instantiation thereof. To initialize a new process instance (e.g., number 476948), participants send their data to the Process Interface (1) to authenticate themselves as participants in the process with their role (e.g., that user 0xc5…Fd43 operates as a WARD clerk in a new process instance PID476948). (2) The authentication is attested to by certifiers confirming the user’s self-declared roles.
@@
When all process participants are identified and associated with a role, (3) the Process Interface merges the received information with the process data acquired in the configuration. Thereupon, (4) a selected participant kickstarts a new process instance through the Process Interface associated with the aforementioned information on the Process Contract, thereby initializing the instance’s public state. This step is done by sending a transaction to the Process Contract with the public state of the process, including conditions driving the control flow, such as exclusive gateway choices (as per Table 1). Notice that this crucial step is notarized on chain (5) (meeting Table 1). The selection of the participant, or participants, on behalf of whom the transaction is signed, transcends the scope of this paper and depends on the decision process and guidelines of the consortium.
@@
After initializing the process state, (6) the Process Interface triggers the Confidentiality Interface, which (7) stores the policy locator previously saved on the Confidentiality Contract. Finally, (8) the Confidentiality Interface notarizes the operation.
@@
Figure 5 displays the transact functionality, which represents the exchange of information during the execution of an activity in the process instance. It encompasses a public and a confidential case, depending on the process participants and their willingness to classify information as disclosable or not. In both cases, this functionality updates the public state in the blockchain according to previous enforcement as per Table 1 (e.g., activity to execute, expected participant) and manages accessibility (meeting Table 1).
Let us begin with the exchange of public data (and the update of the public state). The main steps are highlighted in blue in Fig. 5 and salient on-chain operations are reported in Algorithm 1.
@@
Initially, a user sends the execution data to the Process Interface in the context of an activity (Table 1).
For example, the WARD clerk identified by 0xc5…Fd43 replies to the
“Check availability” activity request by the RADIOLOGY user within process instance 476948.
@@
To do so, the clerk sends a “Check availability” response message wherein the accepted variable is assigned with true since the appointment is confirmed alongside an available date. The boolean answer (accepted) is public data that is used later to make control-flow decisions and advance the state of the process, defining the next activity to perform. In our scenario, the exclusive choice gateway is based on the value of the accepted variable. Then, the Process Interface invokes the Process Contract to request a state update, sending the execution data (Line 1).
@@
In turn, the Process Contract enforces the business and control-flow logic through the run of its routine (based on previously agreed upon and immutable code, see configure functionality), which checks (Line 1) if the executed activity is enabled and if the message sender (in this case, user 0xc5…Fd43) is the expected one (a WARD), finally updating the state (Line 1) if the verification outcome is positive.
@@
This step advances also the state of the process (Line 1), updating the expected next element (here, “Confirm appointment”) and participant (RADIOLOGY). In case the element is a gateway, its logic is automatically executed to advance the control flow of the process (Line 1). In the end, this piece of information is notarized into the Distributed Ledger (meeting Table 1).
@@
For the confidential case, we refer again to Fig. 5 (see the marks in red). Overall, (1) once the Process Interface has received the data, (2) it forwards it to the Confidentiality Interface. The latter (3) retrieves the policy from the Distributed FS, and (4) uses it to encrypt the data via MA-ABE (as per Table 1). For example, considering Fig. 1 and Table 2, the medical prescription of the patient is encrypted with the following policy: {MINISTRY-INSPECTOR or (PID476948 and (PATIENT or RADIOLOGY))}.
@@
In particular, the Confidentiality Interface retrieves the specific authorization for the data to be stored and embeds it in the encryption mechanism, ensuring that only authorized participants can access it. This step is needed only in confidential cases, as in public ones, information is directly stored on the blockchain and intentionally kept publicly readable.
@@
Once encrypted, (5) the Confidentiality Interface stores data in the Distributed FS, (6) its locator in the Confidentiality Contract, and (7) notarizes the operation in the Distributed Ledger (meeting Table 1 and Table 1).
@@
The execution of this routine by the Confidentiality Contract includes a call to the Process Contract (8) to request a state update enclosing the execution data. Following the example, this data is the activity being executed (“Check availability”), the user that performs it (e.g.,, PATIENT, along with their corresponding address 0x2e…6dd9), and the process instance (e.g.,476948).
Then, (9) the Process Contract checks the current public state to update it. The operations involving the smart contracts (10) are notarized in the Distributed Ledger.
The inspect functionality pertains to the access to message data. The public data can be freely retrieved, while the confidential data is restricted to the sole authorized users included in the access policy provided during the encryption phase (see the configure functionality above) (Table 1). To read public data, a process participant or an auditor may send a request to the Process Interface, which in turn retrieves the necessary information from the Process Contract and sends it back to the requester. This is the case, e.g., for the accepted field of the “Check availability” task. Reading confidential data requires more passages and a preliminary step. All users should request at least once (and at every attribute update) a decryption key, to be used for every message they are going to read. For the key request, the process participant (e.g., a WARD clerk) or an auditor (e.g., a HEALTHCARE-INSPECTOR) ask for a decryption key to the Confidentiality Interface. The latter forwards the request to the Confidentiality Contract, which is thus notarized on the Distributed Ledger. Then, the Confidentiality Interface generates the attribute-based key (or a-b key for short) based on the attributes of the user requesting it (e.g., WARD and PID476948, or HEALTHCARE-INSPECTOR) via MA-ABE, and sends it back. The aforementioned step enables the decryption of confidential data. Once the process participants have received the a-b key, they can use it to decrypt the confidential data. To do so, a user makes a request to the Confidentiality Interface. Then, the Confidentiality Interface retrieves the encrypted data from the Distributed FS and forwards it to the user. At this point, the latter uses the a-b key to decrypt the confidential data via MA-ABE. Only if the user’s attributes satisfy the policy previously used to encrypt the data can the user’s a-b key decrypt the confidential data (and the user inspects it in clear).
@@
We remark that an a-b key stays with the user and lasts for as long as their attributes are not updated (if a user is involved in a new process instance, say 476949, a new attribute gets associated with them: PID476949). The same key will be used to access all messages. For example, consider user 0xB0…1AA1 (who holds the attributes PID476948 and PATIENT).
Per the policies in Table 2, their a-b key can decrypt the medical prescription and the registration, but not the report.
The a-b key of the user 0xc5…Fd43 (PID476948 and WARD) can decrypt the report instead, but not the other two documents.
@@
The key, in other words, is associated with the user, and it is the sole instrument to access a document.
Therefore, the Confidentiality Interface does not need to generate a different decryption key for every document and give the same copy to multiple users, nor create numerous copies of the same document for different decryption keys (one per user), thus saving unnecessary data replications and key distributions, and ensuring integrity (Table 1). Also, as the key stays with the users, and the documents are stored and notarized on tamper-proof external infrastructures (the Distributed FS and Blockchain Platform, respectively), the information remains available even should the CONFETTY system stop functioning (Table 1). These aspects have an impact on security considerations, as we discuss next.
@@ 5 Evaluation
To assess the feasibility of the CONFETTY architecture, we perform a two-step evaluation. First, we analyze the CONFETTY threat model, evaluating it against the proposed architecture. Then, we describe the implemented prototype and analyze its performance. The source code of CONFETTY, alongside the experimental data we use and performance analysis results, are available at https://doi.org/10.5281/zenodo.15482587.
5.1 Threat Model
In this section, we identify potential security threats that could compromise data availability (Table 1) and confidentiality (Table 1) within our solution, and outline how the CONFETTY architecture addresses them. Our threat analysis is based on the STRIDE framework [26], a theoretical model that categorizes threats into six distinct groups: spoofing (i.e., impersonation of a legitimate entity), tampering (i.e., unauthorized modification of data to compromise its integrity), repudiation (i.e., the denial of having performed a particular action), information disclosure (i.e., the unauthorized exposure of sensitive data), denial of service (DoS, i.e., the disruption or degradation of system availability), and elevation of privileges (i.e., the unauthorized acquisition of higher-level access rights). We assume users and components to be honest but curious and trusted, respectively. The possible threats thus reside in the communication between them, subject to potential attacks (see Sect. 4.1). Figure 6 displays a Data Flow Diagram (DFD), in which the orange dotted lines indicate the trust boundaries crossing data exchange points between different perimeters, represented by black arrows connecting different elements. The opening side of each boundary indicates the trusted component in the information exchange.
The CONFETTY Boundary (CB) indicates that CONFETTY is the trusted component in the communication with a user (let it be a certifier, a process participant or owner, or an auditor). In this case, a spoofing attack could result in a user impersonating another. Still, this attack is prevented since we assume that each user is linked to the Blockchain Platform credentials. As for tampering, trust is based on the fairness of the user who sends the data. It cannot be the case for a user to repudiate an action because all of them are recorded on the blockchain with a transaction signed with the user’s private key. An attacker could sniff the data a user sends to CONFETTY. However, we resort to TLS channels for those communications, specifically designed to prevent information leakage [28]. DoS attacks are, in principle, feasible unless we pre-filter the range of potential clients invoking the CONFETTY platform. Indeed, all functionalities foresee an initial data input from users (see, e.g., step 1 in Fig. 5) which malicious nodes could exploit with multiple calls and bulky payloads. However, DoS attacks would lead to a disruption or degradation of the system’s availability. Yet, all data, state, and other process instances’ information can be retrieved from the Blockchain Platform and the Distributed Tamper-proof File Storage to reactivate or replicate the system without any data loss. Information disclosure is prevented because all sensitive information is encrypted via MA-ABE. Elevating a user’s privileges is impossible since attributes and permissions are linked to a blockchain account.
The Blockchain Platform is the trusted component in the Blockchain Boundary (BB). In this case, a spoofing attack could result in a malicious version of CONFETTY impersonating the original software object. This attack is prevented since CONFETTY operates as a gateway for transactions ultimately signed by users via their private keys. Information disclosure can be achieved only on public data or resource locators of sensitive data. However, the latter is stored upon MA-ABE-encryption, making its content unreadable by malicious users. Other possible attacks (DoS, tampering, repudiation, elevation of privileges) are counteracted by the design of blockchain protocols. As for the Distributed Tamper-proof File Storage Boundary (FSB), tampering or DoS attack and elevation of privileges are counteracted by design, too. Spoofing and repudiation attacks are ineffective since the authorship of the provided information is proven by signatures of transactions stored on-chain. Since we resort to MA-ABE encryption, information disclosure is negated since the sensitive data stored is encrypted.
@@ 5.2 Implementation
CONFETTY is deployed on three main tiers. An application server hosts the Confidentiality Interface and the Process Interface, developed in Python 3.6.9 and Java SDK-13, respectively. Users communicate with it via TLS communication channels [28]. We utilize IPFS111IPFS: ipfs.tech/; Sepolia: sepolia.dev; Ganache: archive.trufflesuite.com/ganache; Web3.js: web3js.readthedocs.io/en/v1.10.0/. Accessed: July 16, 2025. as the Distributed Tamper-proof File Storage. We employ Sepolia and Ganache as Programmable Blockchain Platforms with their Ethereum Virtual Machine (EVM) as a test environment and local RPC, respectively, connected via Web3.js1.